# Exploring Representations and Interventions in Time Series Foundation Models

### v3

@@ 1 Introduction
Foundation models have taken significant strides in modeling both textual (Brown et al., 2020) and visual (Dosovitskiy et al., 2020) data, and have made complex language and image processing accessible to non-experts. These models are pre-trained on massive internet-scale datasets and can be used to solve multiple tasks across a variety of domains, with little to no adaptation. Recently, a growing body of work (Garza & Mergenthaler-Canseco, 2023; Goswami et al., 2024; Rasul et al., 2024; Das et al., 2024; Woo et al., 2024; Ansari et al., 2024) has extended the benefits of this paradigm to time series data, a modality prevalent in fields such as finance (Taylor, 2008), healthcare (Goswami et al., 2021), and climate science (Schneider & Dickinson, 1974).
@@
We address each of these knowledge gaps by proposing approaches to systematically analyze, probe, and intervene in TSFMs through their learned representations. To address the first gap, we explore the fundamental question: Do different TSFMs learn and organize knowledge in a similar manner? This analysis is crucial for understanding how knowledge is organized in different models and whether representations are redundant across groups of layers, which can be potentially leveraged to improve inference efficiency. Next, we address the second gap by investigating: What human-interpretable concepts can TSFMs learn, and where do these concepts emerge within the latent representations of the model? Building on our layer-wise representation analysis, we propose new methods to identify and localize these concepts across layers. Finally, we tackle the third gap by asking: Can we intervene in TSFMs during inference to guide their output? Through a novel approach, we demonstrate how to steer learned concepts by adjusting representations across layers, allowing the model to generate concept-informed predictions.
@@ 2 Related Work
Time Series Foundation Models (TSFMs). TSFMs are versatile neural networks pre-trained on vast amounts of time series data, and have shown impressive accuracy even in zero-shot settings. Several TSFMs have been proposed (Garza & Mergenthaler-Canseco, 2023; Liu et al., 2023; Das et al., 2024; Woo et al., 2024; Goswami et al., 2024; Ansari et al., 2024; Ekambaram et al., 2024; Talukder et al., 2024). Most of these are based on variations of the Transformer architecture (Vaswani, 2017), with differences in tokenization strategies, pre-training datasets, and the specific tasks they are designed to address. Consequently, we focus on developing methods to understand Transformer-based models. We specifically analyze three representative models: MOMENT (Goswami et al., 2024), Chronos (Ansari et al., 2024), and MOIRAI (Woo et al., 2024), which are fully open-source and offer distinct approaches to time series tokenization (patch vs. discrete), architecture (encoder-only vs. encoder-decoder), and pre-training objectives (imputation vs. forecasting). Recent advancements in TSFMs have focused mostly on improving their capabilities through scaling, specialization, or improving inference efficiency (Liu et al., 2024a, 2023), as well as by incorporating long and multivariate contexts (Liu et al., 2024b; Żukowska et al., 2024). However, there has been limited research on understanding what a specific TSFMs learn. For example, Goswami et al. (2024) showed that MOMENT learns some human interpretable concepts, while Ansari et al. (2024) investigated the failure modes of Chronos in controlled settings, and Potosnak et al. (2024) explored compositional reasoning within time series models, including TSFM.
@@
Identifying Learned Concepts in Pre-trained Models. Understanding the internal representations learned by pre-trained models has been an active area of research, particularly in the context of LLMs and vision models. Previous studies have explored whether individual neurons or directions in a model’s latent space correspond to specific features or concepts (Dalvi et al., 2019; Goh et al., 2021; Gurnee et al., 2023; Elhage et al., 2022). These investigations often focus on identifying linear representations, where features are encoded as linear combinations of neuron activations. Recent work has also employed various probing techniques to classify and interpret these internal representations, addressing aspects such as truthfulness and model robustness (Azaria & Mitchell, 2023; Zou et al., 2023; Burns et al., 2023; Marks & Tegmark, 2023). While prior work has focused on language and vision models, we show that with simple techniques, we can effectively probe TSFMs to identify and localize concepts.
@@ 3 Methods
We study TSFMs using three complementary methods of analysis and interventions concerning model layer-wise representations. In Section 3.1, we examine learned representations through the lens of similarity, uncovering stored knowledge redundancy in TSFM representations. We leverage this redundancy to prune multiple layers of pre-trained models, thereby improving their efficiency. In Section 3.2, we identify specific concepts learned by TSFMs and localize them to specific parts of model layers and tokens. In Section 3.3, we explore the ability to steer model predictions by intervening in model hidden representations across layers, enabling us to guide model output toward concept-informed predictions.
@@ 3.1 Representation Similarity Analysis and Pruning
Figure 1: For each identified block of layers exhibiting redundant representations (red), we remove the internal layers of the block by zeroing out their weights (blue). For example, if a block consists of five layers, we prune layers 2 through 4, retaining only the first and last layers to reduce representation redundancy while maintaining model integrity. More details on pruning can be found in App. C.
@@
Large TSFMs typically learn redundant representations, which often manifest as block-like structures in heatmaps depicting pairwise similarity between layer activations (Figure 1). We can leverage this redundancy to prune TSFMs, enhancing inference speed while preserving accuracy. We build on prior work (Nguyen et al., 2021) and propose a simple layer pruning strategy, which we call Block-wise Pruning, outlined in Algorithm 1. To preserve the structural integrity of each block, we retain the first and last layers of each block while zeroing out the weights of the intermediate layers. The skip connections within transformer blocks ensure that signals and gradients continue to flow through the rest of the network.
Algorithm 1 Block-wise Pruning
@@
        Zero out the weights of layer $l_{j}$ in model $\mathcal{M}$
@@
  return $\mathcal{M}^{\prime}$
To demonstrate the effectiveness of our proposed pruning strategy, we explore two pruning configurations, one in which we prune all redundant blocks, and the other where prune only a single block. We compare the performance of these pruned models to the original, unpruned TSFMs using standard task-specific accuracy metrics (Mean Squared Error and Mean Absolute Error) and efficiency metrics (inference time in milliseconds and theoretical model size in megabytes). We evaluate these models on widely used imputation (Zhou et al., 2021) and forecasting (Ansari et al., 2024) benchmarks in both zero-shot settings and after linear probing (Goswami et al., 2024).
@@
Refer to caption Refer to caption Refer to caption
Figure 3: Examples of synthetic data generated for concept localization, and concept steering experiments include constant signals with varying trend (left), sinusoidal signals with varying frequency (middle), and compositions of constant signals with varying trends and sinusoidal signals, resulting in sinusoidal signals with varying trends (right).
@@ 3.2 Identifying and Localizing Time Series Concepts
We build on the investigation approach outlined in (Marks & Tegmark, 2023). We say that a feature is linearly represented in a foundation model $\mathcal{M}$ if it is represented as a direction in its latent space. If this feature is linearly represented in $\mathcal{M}$, we also want to identify which layer $l$ in $\mathcal{M}$ learns this concept in the most discriminant way. As a concrete example, consider that we want to check whether $\mathcal{M}$ can distinguish between constant and sinusoidal patterns.
@@
By visualizing the scaled LDA values as shown in Figure 2, one can identify which layers and tokens exhibit the highest degree of separation between classes, offering insights into the network’s internal representations for concept intervention techniques.
@@ 3.3 Concept-Informed Predictions via Model Steering
We explore two alternative modalities of intervention: (1) deriving steering vectors using the mean of hidden activations rather than the median, and (2) steering a single token versus all tokens throughout the model. While our methods are applicable to a wide range of transformer-based foundation models, we focus on two prominent TSFM families for brevity: MOMENT (Goswami et al., 2024) and Chronos (Ansari et al., 2024). Both these models are fully open-source, come in different sizes, yet have fundamentally different design choices. For example, Chronos is based on encoder-decoder transformer (in our work we will investigate encoder part) model which takes discretized time series as input, whereas MOMENT is a multi-task, encoder-only model which takes continuous time series patches as input. Since only the Large variant of MOMENT is publicly available at the time of writing this paper, we supplement our representation analysis results with Moirai (Woo et al., 2024), another popular TSFM which comes in different sizes. More information on hyper-parameters can be found in Appendix E.
@@ 4 Results
We observed consistent improvements in memory efficiency and inference speed over their unpruned counterparts. For example, pruning only Block 3 for MOMENT-Large, resulted in a 11% decrease in estimated model size with a 5% speed up in inference. Furthermore, this pruned model had lower zero-shot imputation MAE for 5 of 9 datasets (ETTh2, ETTm1, ETTm2, Exchange, and Weather) as shown in Tab. 3. Chronos-Large results for zero-shot experiments are reported in Tab. 4. Detailed results on memory usage and speed improvements can be found in Tab. 7. While pruning consistently improved memory efficiency and inference speed compared to unpruned counterparts, performance varied across pruning methods and datasets, with some methods exhibiting considerable degradation. In addition to zero-shot experiments, we conducted experiments where models were fine-tuned post-pruning. For this, we applied MOMENT-Large for forecasting to compare a vanilla (unpruned) model to one with all block redundancies pruned, evaluating the impact of the most aggressive pruning approach. Fine-tuning results in Table 1 show that, notably, the pruned model performed nearly as well as the original, underscoring the potential of our block-wise pruning approach to maintain performance while reducing model complexity. Complete finetuning results are provided in Table 6 in Appendix F.
@@
Figure 6: How does model size influence the patterns of learned representations? The emergence of blocks-like patterns in the Large model appears unpredictable from patterns observed in smaller models. More available in the appendix (Fig. 16).
Refer to caption Refer to caption Refer to caption
(i) Pattern (ii) Trend (iii) Periodicity
Figure 7: Progression of linear separability of concepts at the patch level (y-axis) across different layers (x-axis). Linear separability is measured using the Linear Discriminant Ratio (LDR), derived from model embedding statistics for each predicted class: (i) constant vs. sinusoidal patterns, (ii) increasing vs. decreasing trends, and (iii) high vs. low periodicity. Color gradient indicates separability, with dark blue representing low LDR and yellow indicating high LDR. Certain concepts captured by MOMENT-Large exhibit linear separability, but this separability is not uniform across layers; instead, it emerges at specific points in the model. More available in the appendix (Fig. 18).
@@
Steering: Introduce periodicity (i) and trend (ii, iii) to constant time series
Refer to caption Refer to caption Refer to caption
(i) Periodic (ii) Increasing Trend (iii) Decreasing Trend
Compositional Steering: Introduce trend and periodicity to constant time series
(MOMENT (top), Chronos (bottom))
@@
(i) Trend, $\beta=1.0$ (ii) Trend & periodicity, $\beta=0.5$ (iii) Periodicity, $\beta=0$
Figure 8: Visualization of the MOMENT’s reconstruction and the Chronos’s forecasting predictions (bottom), with concept steering applied in the latent space (blue) and the baseline without concept steering (orange). Concept steering effectively transforms concepts in the latent space, resulting in model predictions that align with the intended concepts introduced. Steering results are illustrated for the following experiments: steering a constant signal input to produce (i) a sinusoidal output, (ii) a constant signal with an increasing(slope $>$ 0), and (iii) decreasing trend (slope $<$ 0). For compositional steering experiments, $\alpha$ controls the strength of sinusoidal and increasing trend concepts. When $\beta=0.5$, models are steered towards a combination of increasing trend and sinusoidal pattern. $\beta=0$ results in steering towards a constant signal with an increasing trend, whereas $\beta=1$ only introduces sinusoidal patterns (iii). Detailed results are available in the Appendix F.
@@
Our concept steering interventions effectively transform the latent space of TSFMs, resulting in model predictions that align with the intended concepts, as demonstrated in Fig. 8. We successfully introduced periodicity and trend concepts to constant time series and demonstrated the ability to combine multiple steering vectors to create more complex patterns. By combining steering vectors representing increasing trends and sinusoidal patterns, we were able to steer model predictions towards a combination of these features. To evaluate the effectiveness of steering in the latent space, we analyzed the impact of our interventions on hidden representation, by projecting in a two-dimensional space using Principal Component Analysis (PCA). We found that steering in the latent space is reflected in these lower-dimensional representations, as illustrated in Fig. 10. Notably, the PCA reduction often captured the concept direction as one of the principal components. This can be attributed to the careful design of our synthetic data generation process.
@@ 5 Discussion
We also explored ways to influence model predictions along conceptual directions using steering matrices. Concept steering has many practical applications, including the ability to correct prediction errors in pre-trained models that may arise from out-of-distribution inference, reflecting the effects of exogenous factors on model predictions or inducing prior knowledge not fully captured during training. Additionally, steering provides a method to reduce computational costs by minimizing the need for fine-tuning. It can also be used to introduce inductive biases from various domains into the model. This is particularly valuable for pre-trained models that may lack exposure to specific concepts due to limited or restricted training data. Such inductive biases can be useful in domains like healthcare, where, e.g., knowledge of the excitatory or inhibitory effects of treatments can guide pre-trained model predictions about whether a patient’s vital signs should increase or decrease in response to specific interventions. Moreover, concept steering can be used for data generation, improving data augmentation techniques and creating more diverse datasets to improve TSFM performance across various tasks. Prior work has already shown that data augmentation techniques can improve TSFM generalization by enhancing model robustness and exposing it to a wider variety of patterns (Ansari et al., 2024).
Finally, our findings also underscore the importance of synthetic data in studying and steering TSFMs. As opposed to meticulously curated datasets used to probe large language and vision models, we demonstrate that for time series models, synthetic data generated using simple mechanisms can effectively identify, localize, and probe the linear concepts learned by these models.
@@
This paper provides insights into how a few basic patterns are linearly represented in time series foundation models. Future work must evaluate whether time series foundation models can learn more complex patterns present in real-world time series and whether steering matrices estimated using synthetic data can be used to steer predictions of out-of-distribution, real-world time series. While our methods are broadly applicable to different transformer-based foundation models, future research should explore other architectures such as state space models (Gu & Dao, 2023) and stacked multi-layer perceptrons (Ekambaram et al., 2024). Moreover, future studies should evaluate whether our findings hold for other time series foundation models and tasks such as anomaly detection, and classification. Beyond time series, we hope that our work inspires the use of synthetic data to steer large language and vision models as well.

### v4

@@ 1 Introduction
Foundation models have taken significant strides in modeling both textual (Brown et al., 2020) and visual (Dosovitskiy et al., 2020) data, and have made complex language and image processing accessible to non-experts. These models are pretrained on massive internet-scale datasets and can be used to solve multiple tasks across a variety of domains, with little to no adaptation. Recently, a growing body of work (Garza & Mergenthaler-Canseco, 2023; Goswami et al., 2024; Rasul et al., 2024; Das et al., 2024; Woo et al., 2024; Ansari et al., 2024) has extended the benefits of this paradigm to time series data, a modality prevalent in fields such as finance (Taylor, 2008), healthcare (Goswami et al., 2021), and climate science (Schneider & Dickinson, 1974).
@@
We address each of these knowledge gaps by proposing approaches to systematically analyze, probe, and intervene in TSFMs through their learned representations. To address the first gap, we explore the fundamental question: Do different TSFMs learn and organize knowledge in a similar manner? This analysis is crucial for understanding how knowledge is organized in different models and whether representations are redundant across groups of layers, which can be potentially leveraged to improve inference efficiency. Next, we address the second gap by investigating: What human-interpretable concepts can TSFMs learn, and where do these concepts emerge within the latent representations of the model? Building on our layer-wise representation analysis, we propose new methods to identify and localize these concepts across layers. Finally, we tackle the third gap by asking: Is it possible to intervene in TSFMs during inference by applying conceptual priors to steer time series predictions? Through a novel approach, we demonstrate how to steer learned concepts by adjusting representations across layers, allowing the model to generate concept-informed predictions.
@@ 2 Related Work
Time Series Foundation Models (TSFMs). TSFMs are versatile neural networks pretrained on vast amounts of time series data, and have shown impressive accuracy even in zero-shot settings. Several TSFMs have been proposed (Garza & Mergenthaler-Canseco, 2023; Liu et al., 2023; Das et al., 2024; Woo et al., 2024; Goswami et al., 2024; Ansari et al., 2024; Ekambaram et al., 2024; Talukder et al., 2024). Most of these are based on variations of the Transformer architecture (Vaswani et al., 2017), with differences in tokenization strategies, pretraining datasets, and the specific tasks they are designed to address. Consequently, we focus on developing methods to understand Transformer-based models. We specifically analyze three representative models: MOMENT (Goswami et al., 2024), Chronos (Ansari et al., 2024), and MOIRAI (Woo et al., 2024), which are fully open-source and offer distinct approaches to time series tokenization (patch vs. discrete), architecture (encoder-only vs. encoder-decoder), and pretraining objectives (imputation vs. forecasting). Recent advancements in TSFMs have focused mostly on improving their capabilities through scaling, specialization, or improving inference efficiency (Liu et al., 2024, 2023), as well as by incorporating long and multivariate contexts (Liu et al., 2025; Żukowska et al., 2024). However, there has been limited research on understanding what a specific TSFMs learn. For example, Goswami et al. (2024) showed that MOMENT learns some human interpretable concepts, while Ansari et al. (2024) investigated the failure modes of Chronos in controlled settings, and Potosnak et al. (2024) explored compositional reasoning within time series models, including TSFM.
@@
Identifying Learned Concepts in Pretrained Models. Understanding the internal representations learned by pretrained models has been an active area of research, particularly in the context of LLMs and vision models. Previous studies have explored whether individual neurons or directions in a model’s latent space correspond to specific features or concepts (Dalvi et al., 2019; Goh et al., 2021; Gurnee et al., 2023; Elhage et al., 2022). These investigations often focus on identifying linear representations, where features are encoded as linear combinations of neuron activations. Recent work has also employed various probing techniques to classify and interpret these internal representations, addressing aspects such as truthfulness and model robustness (Azaria & Mitchell, 2023; Zou et al., 2023; Burns et al., 2023; Marks & Tegmark, 2024). While prior work has focused on language and vision models, we show that with simple techniques, we can effectively probe TSFMs to identify and localize concepts.
@@ 3 Methods
We study TSFMs using three complementary methods of analysis and interventions concerning model layer-wise representations. In Section 3.1, we examine learned representations through the lens of similarity, uncovering stored knowledge redundancy in TSFM representations. We leverage this redundancy to prune multiple layers of pretrained models, thereby improving their efficiency. In Section 3.2, we identify specific concepts learned by TSFMs and localize them to specific parts of model layers and tokens. In Section 3.3, we explore the ability to steer model predictions by intervening in model hidden representations across layers, enabling us to guide model output toward concept-informed predictions.
@@ 3.1 Representation Similarity Analysis and Pruning
Figure 1: For each identified block of layers exhibiting redundant representations (red), we skip computation of their internal layers (blue). For example, if a block consists of five layers, we skip layers 2 through 4, retaining only the first and last layers to reduce representation redundancy while maintaining model integrity. More details on pruning can be found in App. C.
@@
Large TSFMs typically learn redundant representations, which often manifest as block-like structures in heatmaps depicting pairwise similarity between layer activations (Figure 1). We leverage this redundancy to prune TSFMs, enhancing inference speed while preserving accuracy. Building on prior work (Nguyen et al., 2021), we propose a simple and effective layer pruning strategy, which we call Block-wise Pruning, outlined in Algorithm 1. To preserve the structural integrity of each block, we retain the first and last layers of each block and skip computation by removing these blocks altogether.
Algorithm 1 Block-wise Pruning (Skipping Computation)
@@
        Remove layer $l_{j}$ from model $\mathcal{M}$
@@
  return pruned model $\mathcal{M}^{\prime}$
To demonstrate the effectiveness of our proposed pruning strategy, we explore two pruning configurations, one in which we prune all redundant blocks, and the other where we prune only a single block. We compare the performance of these pruned models to the original, unpruned TSFMs using standard task-specific accuracy metrics (Mean Squared Error and Mean Absolute Error) and efficiency metrics (inference time in milliseconds and theoretical model size in megabytes). We evaluate these models on widely used imputation (Zhou et al., 2021) and forecasting (Ansari et al., 2024) benchmarks in both zero-shot settings and after linear probing (Goswami et al., 2024). While prior work such as (Nguyen et al., 2021) primarily focused on pruning individual layers, our approach differs by targeting entire blocks of self-similar layers, preserving boundary layers to maintain representational continuity. Furthermore, we extend block-level pruning to large pretrained Time Series Foundation Models, demonstrating practical effectiveness on diverse real-world tasks beyond classification, achieving substantial inference speedups (up to 52%) with minimal performance degradation, even in zero-shot scenarios.
@@
Figure 3: Examples of synthetic data generated for concept localization, and concept steering experiments include constant signals with varying trend (top), sinusoidal signals with varying frequency (middle), and compositions of constant signals with varying trends and sinusoidal signals, resulting in sinusoidal signals with varying trends (bottom).
@@ 3.2 Identifying and Localizing Time Series Concepts
We build on the investigation approach outlined in (Marks & Tegmark, 2024). We say that a feature is linearly represented in a foundation model $\mathcal{M}$ if it is represented as a direction in its latent space. If this feature is linearly represented in $\mathcal{M}$, we also want to identify which layer $l$ in $\mathcal{M}$ learns this concept in the most discriminant way. As a concrete example, consider that we want to check whether $\mathcal{M}$ can distinguish between constant and sinusoidal patterns.
@@
By visualizing the scaled LDR values as shown in Figure 2, one can identify which layers and tokens exhibit the highest degree of separation between classes, offering insights into the network’s internal representations for concept intervention techniques.
@@ 3.3 Concept-Informed Predictions via Model Steering
We explore two intervention techniques: (1) deriving steering vectors using the mean of hidden activations rather than the median, and (2) steering a single token versus all tokens throughout the model. While our methods are applicable to a wide range of transformer-based foundation models, we focus on two prominent TSFM families for brevity: MOMENT (Goswami et al., 2024) and Chronos (Ansari et al., 2024). Both these models are fully open-source, come in different sizes, yet have fundamentally different design choices. For example, Chronos is based on encoder-decoder transformer (in our work we will investigate encoder part) model which takes discretized time series as input, whereas MOMENT is a multi-task, encoder-only model which takes continuous time series patches as input. We supplement our representation analysis results with Moirai (Woo et al., 2024), another popular TSFM which comes in different sizes. More information on hyper-parameters can be found in Appendix E.
@@ 4 Results
We observed consistent improvements in memory efficiency and inference speed compared to unpruned model counterparts. For example, pruning only Block 3 for MOMENT-Large, resulted in a 11% decrease in estimated model size with a 5% decrease in inference time. Furthermore, this pruned model had lower zero-shot imputation MAE for 5 of 9 datasets (ETTh2, ETTm1, ETTm2, Exchange, and Weather) as shown in Tab. 3. Chronos-Large results for zero-shot experiments are reported in Tab. 4. Detailed results on memory usage and speed improvements can be found in Tab. 7. Although pruning consistently improved memory efficiency and inference speed compared to unpruned counterparts, performance varied between pruning methods and datasets, and some methods showed considerable degradation. In addition to zero-shot experiments, we conducted fine-tuning experiments on pruned models—including MOMENT-Large with all block redundancies removed—to evaluate the impact of the most aggressive pruning approach on forecasting performance. Notably, the pruned model performed on par with its unpruned baseline, as shown in Table 1, demonstrating the potential of our block-wise pruning approach to maintain performance while reducing model complexity. Complete finetuning results are provided in Table 6 in Appendix G.
@@
Figure 6: Does model size influence the patterns of learned representations? The emergence of block-like patterns in larger models cannot be directly inferred from the patterns observed in smaller models. (expanded view is Fig. 16).
Figure 7: Linear separability of concepts at the patch level (y-axis) across model layers (x-axis). Yellow indicates higher separability than blue in the colorbar. Linear separability is measured using the Linear Discriminant Ratio (LDR), derived from model embedding statistics for each predicted class (constant vs. sinusoidal patterns). Certain concepts captured by MOMENT-Large are linearly separable, but this separability is not uniform; it emerges at specific layers and patches of the model. More examples of linear separability of concepts are shown in Fig. 14.
@@
Concept Steering: Use conceptual priors to shape time series predictions.
@@
(i) Periodicity, $\beta=0$ (ii) Trend & periodicity, $\beta=0.5$ (iii) Trend, $\beta=1.0$
Figure 8: Visualization of MOMENT reconstruction and Chronos forecasting predictions (bottom), comparing concept steering in the latent space (blue) with the unsteered baseline (orange). The original constant signal and steered output are referred to as Non-perturbed and Perturbed in the figure legend, respectively. A constant time series is first provided as input to both the proposed steering approach and the unsteered baseline. The non-perturbed output remains a constant signal, as expected, while the perturbed output reflects the new concept introduced via the steering matrix—such as trend, seasonality, or both—depending on the specified $\beta$ parameter value. For both outputs, we show the raw model predictions (lighter color) and their moving averages (darker color) to reduce noise artifacts. Steering results are shown for the following experiments: (i) steering a constant signal to have periodicity ($\beta=0$), (ii) steering a constant signal to have an increasing slope, or trend ($\beta=1.0$), and (iii) steering a constant signal to have trend and periodicity ($\beta=0.5$).
@@
Our concept steering interventions effectively transform the latent space of TSFMs, resulting in model predictions that align with the intended concepts, as demonstrated in Fig. 8. We successfully introduced periodicity and trend concepts to constant time series and demonstrated the ability to combine multiple steering vectors to create more complex patterns. By combining steering vectors representing increasing trends and sinusoidal patterns, we were able to steer model predictions towards a combination of these features. To assess the robustness of concept steering, we generated datasets using multiple random seeds and confirmed that consistent steering effects emerge across samples. This is supported by analyses of linear concept separability and concept emergence in the latent space. Furthermore, in Appendix F, we present steering results on real-world ECG data, where we steer time series from the normal to abnormal heartbeat class in the ECG5000 dataset. The ability of our proposed concept steering method to generalize across both synthetic and real-world data highlights its robustness and effectiveness.
To assess the impact of steering in the latent space, we analyzed changes in the hidden representations before and after applying concept steering by projecting them into a two-dimensional space using Principal Component Analysis (PCA). We found that steering in the latent space is reflected in these lower-dimensional representations, as illustrated in Fig. 11. The PCA reduction often captured the concept direction as one of the principal components. This could be attributed to the distinct concepts in our synthetic data.
@@ 5 Discussion
We also explored ways to use conceptual priors to shape time series predictions. Concept steering enables controlled, post-training updates to model embeddings through a lightweight steering matrix, allowing models to incorporate new concepts or events into predictions without requiring retraining or fine-tuning. In addition to reducing the computational cost of further training, concept steering has many practical applications. It enables users to imbue models with new or missing contextual factors after training, which is particularly valuable when models weren’t originally trained on certain scenarios. For example, we can steer vital signs in healthcare based on new treatments or adjust financial forecasts based on emerging events such as positive earnings surprises. Adding these contextual factors, even as simple trends, has significant implications for improving model predictions in zero-shot and out-of-distribution scenarios. Steering also supports synthetic data generation of realistic time series variations. In our experiments with the ECG Arrhythmia classification dataset (ECG5000), we demonstrate that time series classified as normal heartbeats can be steered to produce time series classified as abnormal heartbeats. Such data generation capabilities can be used to augment data for model training or generate new samples for us to better understand the decision boundaries in model predictions. Prior work has already shown that data augmentation techniques can improve TSFM generalization by enhancing model robustness and exposing it to a wider variety of patterns (Ansari et al., 2024).
@@
While our methods are broadly applicable to different transformer-based foundation models, future research should explore other architectures, such as state space models (Gu & Dao, 2024) and stacked multi-layer perceptrons (Ekambaram et al., 2024). Moreover, future studies could evaluate steering methods on other real-world time series datasets and domains, as well as on tasks beyond forecasting and imputation, such as anomaly detection and classification.
Impact Statement
While our pruned TSFMs demonstrate promising performance, it is crucial to exercise caution when using them, especially in high-stakes applications such as healthcare. Before deploying these models for critical decision-making, we strongly recommend fine-tuning and evaluating them on task-specific, in-domain datasets using relevant metrics. The ability to steer model predictions offers numerous benefits but also raises concerns regarding potential biases. We urge users to exercise caution when using our proposed steering strategies and to carefully consider the potential implications of manipulating model outputs.
Reproducibility Statement
To ensure reproducibility, we have made our code anonymously accessible through GitHub. All time series foundation models used for our analysis are publicly available and open-source: Chronos, MOMENT, and MOIRAI. All models were trained and evaluated on a computing cluster consisting of 128 AMD EPYC 7502 CPUs, 503 GB of RAM, and 8 NVIDIA RTX A6000 GPUs, each with 49 GiB RAM.